OpenStudy (anonymous):

Can someone please help me understand this: Find the exact value by using a half-angle identity. sin 7PI/8

OpenStudy (anonymous):

http://www.purplemath.com/modules/idents.htm Here is the half angle id. Just plug and go.

OpenStudy (anonymous):

and your angle is 45 :P

OpenStudy (anonymous):

@Catch.me how do you know it's 45

OpenStudy (anonymous):

did you do 7pi/8 *(pi/180) ?

OpenStudy (anonymous):

right 7 * 180 / 8 multiply it by 2 =7*180/4=315 and it is in the fourth quad ,so its cosine is the same as 45

OpenStudy (anonymous):

Can I set it up like this:

OpenStudy (anonymous):